OpenStudy (anonymous):

if 7x+3=24, what is the value of 5-4x? A. -23 B. -7 C. 1 D.17

OpenStudy (anonymous):

7x+3=24 -3 -3 7x = 21 /7 /7 x=3 5-4(3) = -7

OpenStudy (anonymous):

thanks